The width of a rectangle is 2 ft less than the length. The area is 8ft^2. Find the length and the width.
Found 2 solutions by edjones, RAY100:
Answer by edjones(8007)

L=w+2
wL=8
w(w+2)=8
w^2+2w-8=0
(w+4)(w-2)=0
w=2'
L=4'
.
Ed

Answer by RAY100(1637)

Area = Length * Width
.
but W=L-2
.
A=LW = L(L-2) = L^2 -2L = 8,,,,,given
.
L^2 -2L -8 =0,,,,,,,,,factor
.
(L-4)(L+2)=0
.
L= 4,,-2 ( not realistic)
.
W= L-2 = 4-2 = 2
.
check,,,,A=LW=4*2=8,,,,,ok